ISS Daily Summary Report – 10/15/15

Plant Gravity Sensing 2 (PGS2) Run 1:  Following up on Lindgren’s treatment yesterday of the growing plants with a chemical reagent that induces a bioluminescence response, today Yui removed the reagent with an absorbent and placed them in a Photon Counting Unit (PCU).  Photon measurements will be taken over the next 24 hours.  Lindgren noted yesterday that one of the four culture dishes was contaminated and, after coordination with ground experts, he discarded that dish.  Plant calcium concentrations have been shown to change in response to the direction of gravity:  the treated plants will emit photons when the culture dishes they are growing in are rotated so that they are “upside down.”  The PGS2 investigation supports the study of cellular formation of the plant’s gravity sensors and the molecular mechanism for gravity sensing in plants grown in microgravity conditions.

 

MAGVECTOR:  Today Lindgren completed the sixth run of the European Space Agency’s (ESA’s) MAGVECTOR experiment begun last Friday, October 9th, by transferring data from a jump drive to a Station Support Computer (SSC).  MAGVECTOR qualitatively investigates the interaction between a moving magnetic field and an electrical conductor. The expected changes in the magnetic field structure on the ram and wake side of the electrical conductor are of interest for technical applications as well as for astrophysical research.

 

SOLAR:  Measurements continue to be taken for European Space Agency’s (ESA’s) SOLAR investigation during the current sun visibility window which is open from October 9th to October 19th.  The goal of the SOLAR instruments is to measure solar spectral irradiance and variability.

 

Journals:  Kelly made Journals entries today. The Journals investigation obtains information on behavioral and human issues that are relevant to the design of equipment and procedures used during astronauts during extended-duration missions. Study results provide information used in preparation for future missions to low-Earth orbit and beyond.

 

Space Headaches: Yui answered his weekly Space Headaches questionnaire today.  Headaches can be a common complaint during spaceflight. The Space Headaches experiment will provide information that may help in the development of methods to alleviate associated symptoms and improvement in the well-being and performance of crew members in space.

 

Extravehicular Activity (EVA) Preparation:  Lindgren and Kelly obtained body measurements on each other as part of Extravehicular Mobility Unit (EMU) On-Orbit Fit Verification (OFV).  The measurements will be used to confirm the correct sizing of the EMU suits prior to going EVA on October 28th.  In addition, they installed Rechargeable EVA Battery Assemblies (REBA) on EMU suits 3003 and 3010.  Once installed, they performed a checkout to verify the EMU Glove heaters were functional and helmet cameras were receiving power.   Lastly, Lindgren replaced the gas trap on EMU 3010.

 

Ocular Health:  Yui and Lindgren executed their Medical Operations Flight Day 90 (FD90) Ocular Health activities with Kelly’s assistance.  Both performed Optical Coherence Tomography (OCT) and fundoscope measurements with Kelly as operator.

 

Today’s Planned Activities
All activities were completed unless otherwise noted.

  • HMS – OCT Hardware Setup
  • Terminate EMU LIB Battery Charge Operations
  • SEISMOPROGNOZ. Downlink data from Control and Data Acquisition Module (МКСД) HDD (start)
  • Installation of СУБА Hardware (as part of КЛ-121/122Ц monoblock installation)
  • Private Psychological Conference
  • Start EMU Battery Charge
  • WRS Water Sample Analysis
  • Water Recovery System (WRS). WPA Waste Water Tank Offload into CWC (start)
  • PILOT-T. Experiment Ops
  • PGS2 – Equipment Gathering
  • EHS MCD – In-flight Microbiology Water analysis and data recording
  • Water Recovery System (WRS). Terminate WPA Waste Water Tank Offload into CWC
  • СОЖ Maintenance
  • OCT Eye Examination, Operator
  • OCT Eye Examination, Subject
  • Ultrasound 2 HRF – Rack 1 Power Up
  • Ultrasound – Scan Prep
  • OCT Eye Examination, Operator
  • OCT Eye Examination, Subject
  • SEISMOPROGNOZ. Downlink data from Control and Data Acquisition Module (МКСД) HDD (end) and start backup
  • Eye Examination (Operator)
  • Study of cardiovascular system under graded physical load on VELO (Operator)
  • Study of Cardiovascular System Under Graded Physical Load on VELO
  • Eye Examination (Subject) / See OPTIMIS Viewer for Procedure
  • Ultrasound 2 – Scan by CMO
  • Ultrasound 2 – Scan (Subject) / See OPTIMIS Viewer for Procedure
  • OCT Hardware Stowage
  • TOCA Data Recording
  • XF305– Camcorder Setup
  • Plant Gravity Sensing (PGS2). Absorbing chemical solution
  • Ultrasound 2 – Closeout Ops
  • Ultrasound Data Export
  • Acoustic Dosimeter Ops
  • КЦП1 and Laptop RS1 Software Update
  • Ultrasound 2 – Deactivation and Stowage
  • Photography of ISS RS window 06, 07, 08, 09, 26 glass
  • R/G review, search and gathering of required equipment prior to replacement of ПВ-12Р plug
  • MAGVEC – Data Export
  • Replacement of fuses in ПВ-12Р plug
  • Journal Entry
  • HMS – Food Frequency Questionnaire
  • Water sample collection from SVO-ZV (EDV 1198)
  • SHD – Weekly Questionnaire
  • EMU On-Orbit Fit Check
  • Node 1 Nadir – Connecting Power Cables Part 3
  • EMU On-Orbit Fit Check
  • Replacement of damaged plug ПВ-12(#10Ю=А811) on БСКУ5-12 (#10ЮА781) with ПВ-12Р plug taken from onboard spares
  • Installation of REBA batteries
  • REBA powered equipment test
  • CALCIUM. Experiment Session 9
  • MOTOCARD. Experiment Ops
  • MOTOCARD. Operator Assistance with the Experiment
  • EMU Gas Trap Remove and Replace
  • EXPANDER Exercise
  • ARED Exercise Video Equipment Stowage
  • Fundoscope Hardware Setup
  • Fundoscope – Eye Exam (Operator)
  • Fundoscope Hardware Setup
  • Fundoscope – Eye Examination
  • Fundoscope Power Down and Stowage
  • Preparation of reports for Roscosmos site
  • ECON-M. Observations and Photography
  • IMS Delta File Prep

 

Completed Task List Items

  • P/TV N1 Video Cable R&R
  • 61P USOS cargo unpack [In work]

 

Ground Activities

All activities were completed unless otherwise noted.

  • HCZ MDM Swap and EEPROM refresh

 

Three-Day Look Ahead:

Friday, 10/16:  Fundoscope, Ocular Health, X2R14, IMAX, JEMAL Shielding, EVA LLB Battery Charge

Saturday, 10/17:  Weekly Cleaning, Crew Off Duty

Sunday, 10/18: Crew Off Duty

 

QUICK ISS Status – Environmental Control Group:

                              Component Status
Elektron On
Vozdukh Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) On
Carbon Dioxide Removal Assembly (CDRA) Lab Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 Operate
Major Constituent Analyzer (MCA) Lab Shutdown
Major Constituent Analyzer (MCA) Node 3 Operate
Oxygen Generation Assembly (OGA) Process
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full Up